### v3

# Deep Reinforcement Learning for Robotic Bipedal Locomotion: A Brief Survey

Abstract
Bipedal robots are gaining global recognition due to their potential applications and advancements in artificial intelligence, particularly through Deep Reinforcement Learning (DRL). While DRL has significantly advanced bipedal locomotion, the development of a unified framework capable of handling a wide range of tasks remains an ongoing challenge. This survey systematically categorises, compares, and analyses existing DRL frameworks for bipedal locomotion, organising them into end-to-end and hierarchical control schemes. End-to-end frameworks are evaluated based on their learning approaches, while hierarchical frameworks are examined in terms of layered structures that integrate learning-based or traditional model-based methods. We provide a detailed evaluation of the composition, strengths, limitations, and capabilities of each framework. Additionally, this survey identifies key research gaps and proposes future directions aimed at creating a more integrated and efficient unified framework for bipedal locomotion, with wide-ranging applications in real-world environments.
Index Terms:
Deep Reinforcement Learning, Humanoid Robots, Bipedal Locomotion, Legged Robots
I Introduction
Humans navigate complex environments and perform diverse locomotion tasks with remarkable efficiency using just two legs. Bipedal robots, which closely mimic human form, have a distinct advantage over wheeled or tracked alternatives, particularly when traversing uneven and challenging terrains. Furthermore, bipedal humanoid robots are specifically designed to function in human-centric environments, enabling seamless interaction with tools and infrastructure intended for human use. This makes them highly adaptable to a wide range of tasks in such settings.
As a result, bipedal robots hold significant potential for real-world applications [1]. In manufacturing, they can perform tasks efficiently without requiring additional tools, thereby enhancing productivity and reducing labour demands [2, 3, 4]. Their agility is particularly advantageous in complex environments, such as multi-level workplaces. Bipedal robots are also well-suited for tasks that involve the use of human-designed tools, making them valuable for assisting in daily tasks, healthcare, and rehabilitation [5]. Furthermore, they show considerable promise in search and rescue operations, where they can navigate hazardous and unpredictable terrains [6, 7, 8].
Traditional approaches to bipedal locomotion control, such as model-based methods, have been prevalent since the 1980s [9, 10, 11]. Early methods, such as the Linear Inverted Pendulum Model (LIPM) [12], allowed for simplified representations of the dynamics involved in bipedal movement, enabling easier analysis and control. As research progressed, full dynamic models were introduced to better capture the complexities of real-world locomotion. While model-based approaches offer rapid convergence and predictive capabilities, they often fall short in dynamically complex and uncertain environments, where adaptability is key.
Reinforcement Learning (RL)-based methods, particularly Deep Reinforcement Learning (DRL), are effective in optimising robot control policies through direct interaction with the environment [13], which offers a distinct advantage. Unlike model-based approaches, which rely on predefined dynamics and can struggle with unforeseen scenarios, DRL enables robots to autonomously discover control strategies through trial and error, allowing for greater adaptability and robustness in diverse environments. Additionally, hybrid methods that combine model-based and learning-based techniques further enhance planning and control by leveraging the strengths of both approaches.
Figure 1: Common bipedal and humanoid robots used as platforms for testing DRL frameworks. (a) NAO, a toy-like 3D humanoid robot, actuated by servo motors [14]. (b) Rabbit, a 2D bipedal robot, actuated by torque control [15]. (c) Cassie, a 3D bipedal robot, also actuated by torque control [16]. (d) ATLAS, a 3D humanoid robot, driven by hydraulics [17]. (e) Digit, a full human-sized 3D humanoid robot, an upgrade based on Cassie and actuated by torque control [18].
Despite these advancements, a key question remains: Is there a unified framework capable of enabling bipedal robots to effectively manage a diverse range of locomotion tasks? If not, how should such a framework be designed, and what features or components should it include?
To address these questions, this survey examines recent advancements in DRL-based frameworks, categorising control schemes into two primary types: (i) end-to-end and (ii) hierarchical. These frameworks enable robots to generate diverse gaits, including walking, running, and turning, while also allowing them to navigate complex terrains, such as stairs, slopes, and deformable surfaces.
End-to-end frameworks directly map robot states to joint-level control outputs, while hierarchical frameworks decompose decision-making into multiple layers. In hierarchical systems, a High-Level (HL) planner governs navigation and path planning, while a Low-Level (LL) controller handles fundamental locomotion tasks. The task-level decision-making tier interfaces directly with user commands or predefined tasks, forming a structured approach to robotic control.
The evolution of RL in bipedal robotics has spurred a dynamic growth in innovative applications. Although the application of RL to simple 2D bipedal robots began in 2004 [19, 20], it took several years before DRL algorithms emerged. These DRL-based methods have since shown promising results in physical simulators [21, 22, 23]. Agility Robotics introduced the first sim-to-real end-to-end learning framework in 2019, which was applied on the 3D torque-controlled bipedal robot Cassie, as shown in Fig. 1(c) [16]. Besides model-based reference learning, the policy can also incorporate motion capture data [24, 25, 26, 27], or start from scratch [28] to explore solutions freely. Recent studies demonstrate that end-to-end frameworks robustly handle complex and diverse tasks [29, 30, 31].
Similarly, hierarchical structures have garnered significant interest. Within this subset, the hybrid approach combines RL-based and model-based methods to enhance both planning and control strategies. A notable framework employs a learned HL planner coupled with a LL model-based controller, often referred to as the Cascade-structure or Deep Planning Hybrid Scheme [32, 33, 34]. Another innovative construction integrates a learned feedback controller with an HL planner, categorised under the DRL Feedback Control Hybrid Scheme [35, 36]. Additionally, a learned hierarchical control scheme [37] decomposes locomotion into various tasks, focusing each layer on specific functions such as navigation and fundamental locomotion skills [21, 22, 38]. In order to provide a clearer understanding of the current landscape, we have categorised existing DRL frameworks, as depicted in Fig. 2.
Current progress shows that we are still far from achieving a unified framework, despite two mainstream types of framework structure. Although bipedal locomotion tasks have evolved significantly in the past decade. Earlier frameworks primarily focused on basic walking stabilization and simple terrain adaptation, whereas recent developments have introduced complex tasks such as dynamic stair climbing, parkour-inspired locomotion, and loco-manipulation. Notably, the DARPA Robotics Challenge (DRC) has established a structured benchmark for bipedal humanoid locomotion and manipulation [39], emphasizing real-world deployment and robustness.
Although several reviews discuss RL for general robotics [13] and model-based methods for bipedal robots [9, 10, 11], none specifically focus on DRL-based frameworks for bipeds. To address this gap, this survey reviews relevant literature by applying the following selection criteria: (1) studies that examine DRL frameworks specifically for bipedal robots, (2) research involving both simulated and physical bipedal robots, (3) approaches that improve policy transfer from simulation to real-world environments, and (4) publications from the last five years (2018 - April 2024) sourced from reputable databases and conferences, including Google Scholar, IEEE Xplore, Web of Science, ArXiv, and major robotics conferences such as CoRL, RSS, ICRA, IROS, and Humanoids.
The search was conducted using the key terms “Deep Reinforcement Learning” or “Reinforcement Learning” in combination with “Bipedal Locomotion”, “Bipedal Walking”, “Biped Robot”, “Humanoid Robot”, or “Legged Robot”. The most relevant and impactful works were manually selected for further review. This survey is intended for readers with a foundational background in robotics who are transitioning to DRL methods, providing an overview of a wide range of approaches with simplified explanations where necessary. For clarity, in this survey, the term “humanoid robot” refers specifically to anthropomorphic bipedal robots.
The primary contributions of this survey are:
  • A comprehensive summary and cataloguing of DRL-based frameworks for bipedal locomotion.
  • A detailed comparison of each control scheme, highlighting their strengths, limitations, and characteristics.
  • The identification of current challenges and the provision of insightful future research directions.
The survey is organised as follows: Section II focuses on end-to-end frameworks, categorised by learning approaches. Section III details hierarchical frameworks, classified into three main types. Section IV presents the expected unified frameworks, which build upon proposed approaches while highlighting the challenges associated with their implementation. Section V addresses potential future research directions. Finally, Section VI concludes the survey.
Figure 2: Classification of DRL-based control schemes.
II End-to-end framework
The end-to-end DRL framework represents a holistic approach where a single neural network (NN) policy, denoted $\pi(\cdot):\mathcal{X}\rightarrow\mathcal{U}$, directly maps sensory inputs $\mathcal{X}$, such as images, lidar data, or proprioceptive feedback [40], along with user commands [28] or pre-defined references [41], into joint-level control actions $\mathcal{U}$. Here, $\mathcal{X}$ represents the sensory input space, $\mathcal{U}$ refers to the space of control actions, and $\pi(\cdot)$ denotes the policy function. This framework obviates the need for manually decomposing the problem into sub-tasks, streamlining the control process.
End-to-end strategies primarily simplify the design of LL tracking to basic elements, such as a Proportional-Derivative (PD) controller. These methods can be broadly categorised based on their reliance on prior knowledge into two types: reference-based and reference-free. The locomotion skills developed through these diverse learning approaches exhibit considerable variation in performance and adaptability.
In the following sections, we will delve into various representation frameworks, exploring their characteristics, limitations, and strengths in comprehensive detail. To facilitate an understanding of these distinctions, Table I provides a succinct overview of the frameworks discussed.
TABLE I: Summary and comparison of reference-based and reference-free Learning approaches for end-to-end framework. The dashed line in the implementation flow chat refers to optional.
Methods Works Capabilities Characteristic Advantages and Disadvantages Implementation Flow Chart
Residual learning [42] [43] Forward walk undirectional walk Add residual term to the known motor positions at the current time step. A: Fast convergence speed D: Require high-quality predefined reference, limit to specific motions, and lack robustness to complicated terrains. Refer to caption
Guided learning [44] [41] [45] [31] Forward walk Versatile walk Versatile jump Versatile motions Mimic the predefined reference and directly specifies joint-level command. A: Accelerate the learning process and robust to the terrains. D: Limited to the predefined motions and lack adaptability to unforeseen changes in environment. Refer to caption
Reference-free learning [28] [46] [29] Periodic motions Stepping stones walk Visual walk Learn locomotion skills from scratch without any prior knowledge. A: High potential for gait exploration, high robust to the complicated terrain. D: Requires intensive reward shaping for gait patterns and relatively expensive computational training resources. Refer to caption
00footnotetext: Forward Walk involves the bipeds walking straight ahead. Unidirectional Walk enables the bipeds to move both forward and backward within a range of desired velocities. Omni-Walk grants the bipeds the ability to walk in any direction. Versatile Walk allows the bipeds to walk forward, backward, turn, and move sideways, providing extensive movement capabilities. Periodic Motions entails the execution of various repeated gait patterns, such as walking, hopping, or galloping. Versatile Jump refers to jump towards different desired targets. Versatile Motions cover performing a broad array of motions, both periodic and aperiodic such as jumping.
II-A Reference-based learning
Reference-based learning leverages prior knowledge generated offline through methods such as trajectory optimisation (TO) or motion capture systems. This predefined reference typically includes data related to the robot’s joint movements or pre-planned trajectories, serving as a foundation for the policy to develop locomotion skills by following these established motion patterns. Generally, this approach can be divided into two primary methods: (i) residual learning and (ii) guided learning.
II-A1 Residual learning
The proposed framework utilises a policy that modifies motor commands by applying action offsets based on the current reference joint positions, allowing the biped robot to achieve dynamic locomotion through error compensation. The state space includes proprioceptive information such as trunk position, orientation, velocity, angular velocity, joint angles, and joint velocities, providing the necessary sensory data for real-time adjustments. Actions are defined by offsets, $\delta a$, which represent deviations from the predefined desired joint positions, $\hat{a}$, with the final motor commands represented as $a=\hat{a}+\delta a$. The reward function encourages the policy to optimise locomotion performance by considering (a) how closely the robot’s active joint angles match the reference angles, (b) how effectively the robot responds to user commands, and (c) additional terms that further enhance the stability of the robot’s movements. This holistic approach enables the biped robot to adapt to various dynamic conditions while maintaining balance and control.
Introduced in 2018, a residual learning framework for the bipedal robot Cassie marked a significant advancement [42]. This framework allowed the robot to walk forward by incorporating a policy trained via Proximal Policy Optimisation (PPO) algorithms, as detailed in Appendix A. The policy receives the robot’s states and reference inputs, outputting a residual term that augments the reference at the current timestep. These modified references are then processed by a PD controller to set the desired joint positions. While this framework enhanced the robot’s ability to perform tasks beyond standing [47], its physical deployment on a bipedal robot has not yet occurred, potentially rendering it impractical for managing walking at varying speeds and limiting movement to a single direction.
To transition this framework to a real robot, a sim-to-real strategy based on the previous model was demonstrated, where the policy, trained through a residual learning approach, was subsequently applied on a physical bipedal robot [43]. Compared to model-based methods, this training policy achieves faster running speeds on the same platform, underlining the considerable potential of DRL-based frameworks. However, the robot’s movements remain constrained to merely walking forward or backward.
A unique residual learning approach was introduced to enable omnidirectional walking, where the policy adds a residual term to the current joint positions, allowing gradual omnidirectional walking [48]. In this case, the desired reference is the robot’s current joint positions, which makes the approach distinctive. However, this also limits the policy’s ability to explore more diverse motions, restricting it to a single slow walking pattern.
Residual learning enhances an existing control policy by taking current joint positions or states and applying a residual action to adjust reference actions for better performance. Compared to other learning approaches that directly output joint positions, it is highly sample-efficient [33]. However, when predefined references are unstable or of low quality, residual learning may struggle, especially on complex terrains, as the action space is bounded by the reference, limiting the ability to handle unpredictable or uneven terrains.
II-A2 Guided learning
Guided learning trains policies to directly output the desired joint-level commands as actions $a$, without relying on the addition of a residual term. The states space is as same as residual learning approach. In this approach, the reward structure is centred on accurately imitating predefined reference trajectories, ensuring precise alignment between the policy output and the reference motion.
A sim-to-real framework that employs periodic references to initiate the training phase was proposed in [44]. In this framework, the action space directly maps to the joint angles, and desired joint positions are managed by joint PD controllers. The framework also incorporates a Long Short-Term Memory (LSTM) network, as detailed in Appendix A, which is synchronised with periodic time inputs. However, this model is limited to a single locomotion goal: forward walking. A more diverse and robust walking DRL framework that includes a Hybrid Zero Dynamics (HZD) gait library was demonstrated [41], achieving a significant advancement by enabling a single end-to-end policy to facilitate walking, turning, and squatting.
Despite these advancements, the parameterisation of reference motions introduces constraints that limit the flexibility of the learning process and the policy’s response to disturbances. To broaden the capabilities of guided learning policies, a framework capable of handling multiple targets, including jumping, was developed [45]. This approach introduced a novel policy structure that integrates long-term input/output (I/O) encoding, complemented by a multi-stage training methodology that enables the execution of complex jumping maneuvers. An adversarial motion priors approach, employing a style reward mechanism, was also introduced to facilitate the acquisition of user-specified gait behaviours [27]. This method improves the training of high-dimensional simulated agents by replacing complex hand-designed reward functions with more intuitive controls.
While previous works primarily focused on specific locomotion skills, a unified framework that accommodates both periodic and non-periodic motions was further developed [31] based on the foundational work in [45]. This framework enhances the learning process by incorporating a wide range of locomotion skills and introducing a dual I/O history approach, marking a significant breakthrough in creating a robust, versatile, and dynamic end-to-end framework. However, experimental results indicate that the precision of locomotion features, such as velocity tracking, remains suboptimal.
Guided learning methods expedite the learning process by leveraging expert knowledge and demonstrating the capacity to achieve versatile and robust locomotion skills. Through the comprehensive evaluation [31], it is demonstrated that guided learning employs references without complete dependence on them. Conversely, residual learning exhibits failures or severe deviations when predicated on references of inferior quality. This shortfall stems from the framework’s dependency on adhering closely to the provided references, which narrows its learning capabilities.
Nonetheless, reference-based learning reliance on predefined trajectories confines the policy to specific gaits, restricting its capacity to explore a broader range of motion possibilities [41, 49]. Additionally, this approach exhibits limited adaptability in responding effectively to unforeseen environmental changes or novel challenges.
II-B Reference-free learning
In reference-free learning, the policy is trained using a carefully crafted reward function rather than relying on predefined trajectories. This approach allows the policy to explore a wider range of gait patterns and adapt to unforeseen terrains, thereby enhancing innovation and flexibility within the learning process. The action space and observation space in this approach are similar to the guided-learning method; however, the reward structure differs significantly from the reference-based method. Instead of focusing on imitating predefined motions, the reward emphasises learning efficient gait patterns by capturing the distinctive characteristics of bipedal locomotion [49].
The concept of reference-free learning was initially explored using simulated physics engines with somewhat unrealistic bipedal models. A pioneering framework, which focused on learning symmetric gaits from scratch without the use of motion capture data, was developed and validated within a simulated environment [23]. This framework introduced a novel term into the loss function and utilised a curriculum learning strategy to effectively shape gait patterns. Another significant advancement was made in developing a learning method that enabled a robot to navigate stepping stones using curriculum learning, focusing on a physical robot model, Cassie, though this has yet to be validated outside of simulation [50].
Considering the practical implementation of theoretical models, significant efforts have been directed towards developing sim-to-real frameworks in robotics studies. A notable example of such a framework accommodates various periodic motions, including walking, hopping, and galloping [28]. This framework employs periodic rewards to facilitate initial training within simulations before successfully transitioning to a physical robot. It has been further refined to adapt to diverse terrains and scenarios. For instance, robust blind walking on stairs was demonstrated through terrain randomisation techniques in [46]. Additionally, the integration of a vision system has enhanced the framework’s ability to precisely determine foot locations [51], thus enabling the robot to effectively navigate stepping stones [29]. Subsequent developments include the incorporation of a vision system equipped with height maps, leading to an end-to-end framework that more effectively generalises terrain information [52].
This approach to learning enables the exploration of novel solutions and strategies that might not be achievable through mere imitation of existing behaviours. However, the absence of reference guidance can render the learning process costly, time-consuming, and potentially infeasible for certain tasks. Moreover, the success of this method hinges critically on the design of the reward function, which presents significant challenges in specifying tasks such as jumping.
III Hierarchy framework
Unlike end-to-end policies that directly map sensor inputs to motor outputs, hierarchical control schemes deconstruct locomotion challenges into discrete, manageable layers or stages of decision-making. Each layer within this structure is tasked with specific objectives, ranging from navigation to fundamental locomotion skills. This division not only enhances the framework’s flexibility but also simplifies the problem-solving process for each policy.
The architecture of a hierarchical framework typically comprises two principal modules: an High-Level (HL) planner and an Low-Level (LL) controller. This modular approach allows for the substitution of each component with either a model-based method or a learning-based policy, further enhancing adaptability and customisation to specific needs.
Communication between the layers in a hierarchical framework is achieved through the transmission of commands. The HL planner sets abstract goals, which the LL controller translates into specific actions, such as calculating joint movements to follow a desired trajectory. In return, the robot sends sensor data back to the HL planner, enabling real-time adjustments. The tasks handled by different layers often operate on varying time scales, adding complexity to synchronising communication between the layers.
Hierarchical frameworks can be classified into three distinct types based on the integration and function of their components:
  1. 1.
    Deep planning hybrid scheme: This approach combines strategic, HL planning with dynamic LL execution, leveraging the strengths of both learning-based and traditional model-based methods.
  2. 2.
    Feedback DRL control hybrid scheme: It focuses on integrating direct feedback control mechanisms with DRL, allowing for real-time adjustments and enhanced responsiveness.
  3. 3.
    Learned hierarchy scheme: Entirely learning-driven, this scheme develops a layered decision-making hierarchy where each level is trained to optimise specific aspects of locomotion.
These frameworks are illustrated in Fig. 3. Each type offers unique capabilities and exhibits distinct characteristics, albeit with limitations primarily due to the complexities involved in integrating diverse modules and their interactions.
For a concise overview, Table 3 summarises the various frameworks, detailing their respective strengths, limitations, and primary characteristics. The subsequent sections will delve deeper into each of these frameworks, providing a thorough analysis of their operational mechanics and their application in real-world scenarios.
TABLE II: Summary and comparison of Hierarchy framework
Control Scheme Works Module Characteristic Advantages and Disadvantages
Deep Planning Hybrid Scheme [33] [53] [54] Deep planning + ID Deep planning + ID-QP Deep planning + WPG HL policy is learned to guide the LL controller to complete locomotion and navigation tasks. A: Enhanced command tracking capabilities, generalised across different platforms, sampling efficiency, and robust. D: Complicate system and communication between layers, require precise model, lack generalisation regarding different tasks.
Feedback DRL Control Hybrid Scheme [55] [35] [36] Gait library + feedback policy Footstep planner + feedback policy Model-based planner + feedback policy LL feedback policy receives non-learned HL planner as input to achieve locomotion skills. A: Short inference times, robust, navigation locomotion capabilities, interpretability. D: Complicated system and communication between layers, reducing sampling efficiency.
Learned Hierarchy Framework [21] [38] [56] HL policy + LL policy HL policy + LL policy HL policy + LL policy Both HL planner and LL feedback controller are learned. LL policy focuses on basic locomotion skills; on the other side, HL policy learn navigation skills. A: provides layer flexibility, where each layer can be independently retrained and reused; alleviates the challenges associated with training an end-to-end policy. D: inefficienty sim-to-real, complicated interface between layers, training expensively.
III-A Deep planning hybrid scheme
In this scheme, robots are pre-equipped with the ability to execute basic locomotion skills such as walking, typically managed through model-based feedback controllers or interpretable methods. The addition of an HL learned layer focuses on strategic goals or the task space, enhancing locomotion capabilities and equipping the robot with advanced navigation abilities to effectively explore its environment.
Several studies have demonstrated the integration of an HL planner policy with a model-based controller to achieve tasks in world space. A notable framework optimises task space level performance, eschewing direct joint level and balancing considerations [33]. This system combines a residual learning planner with an inverse dynamics controller, enabling precise control over task-space commands to joint-level actions, thereby improving velocity tracking, foot touchdown location, and height control. Further advancements include a hybrid framework that merges HZD-based residual deep planning with model-based regulators to correct errors in learned trajectories, showcasing robustness, training efficiency, and effective velocity tracking [34]. These frameworks have been successfully transferred from simulation to reality and validated on robots such as Cassie.
However, the limitations imposed by residual learning constrained the agents’ capacity to explore a broader array of possibilities. Building on previous work [34], a more efficient hybrid framework was developed, which learns from scratch without reliance on prior knowledge [53]. In this approach, a purely learning-based HL planner interacts with an LL controller using an Inverse Dynamics with Quadratic Programming formulation (ID-QP). This policy adeptly captures dynamic walking gaits through the use of reduced-order states and simplifies the learning trajectory. Demonstrating robustness and training efficiency, this framework has outperformed other models and was successfully generalised across various bipedal platforms, including Digit, Cassie, and RABBIT.
In parallel, several research teams have focused on developing navigation planners specifically for toy-like humanoid robots, which provide greater physical stability compared to torque-driven or hydraulic bipedal robots as shown in Fig. 1. One notable study [57] implemented a visual navigation policy on the NAO robot, depicted in Fig. 1(a), utilising RGB cameras as the primary sensory modality. This system has demonstrated successful zero-shot transfer to real-world scenarios, enabling the robot to adeptly navigate around obstacles. Further research [54] has explored complex dynamic motion tasks, such as playing soccer, by integrating a learned policy with an online footstep planner that utilises weight positioning generation (WPG) to create a centre of mass (CoM) trajectory. This configuration is coupled with a whole-body controller, facilitating dynamic activities like soccer shooting. Despite their platform’s stability, provided by large feet and a lightweight structure, these robots exhibit limited dynamic movement capabilities compared to full-sized humanoid robots. Consequently, this research primarily addresses navigation and task execution.
Regarding generalisation, these frameworks have shown potential for adaptation across different types of bipedal robots with minimal adjustments, demonstrating advanced user command tracking [53] and sophisticated navigation capabilities [54]. However, limitations are evident, notably the absence of capabilities for executing more complex and dynamic motions, such as jumping. Furthermore, while these systems adeptly navigate complex terrains with obstacles, footstep planning alone is insufficient without concurrent enhancements to the robot’s overall locomotion capabilities. Moreover, the requisite communication between the two distinct layers of the hierarchical framework may introduce system complexities. Enhancing both navigation and dynamic locomotion capabilities within the HL planner remains a significant challenge.
Figure 3: Hierarchy Control Scheme Diagram: presents a hierarchical control framework for a bipedal robot, consisting of a basic scheme and three variations. (1) Basic Scheme: The framework starts with a Task Command, followed by a HL Planner, and a LL Controller, which ultimately drives the Robot. Each module can be replaced by a learned policy, introducing adaptability in different control layers. (2) Variations (from left to right): (a) A deep planning hybrid scheme, where the HL planner is learned. (b) A Feedback DRL control hybrid scheme, with a learned LL controller. (c) A Learned hierarchy control scheme, where both layers are learned.
III-B Feedback DRL control hybrid scheme
In contrast to the comprehensive approach of end-to-end policies discussed in Section II, which excels in handling versatile locomotion skills and complex terrains with minimal interface times, the Feedback DRL Control Hybrid Scheme integrates DRL policies as LL controllers. These LL controllers, replacing traditional model-based feedback mechanisms, work in conjunction with HL planners that process terrain information, plan future walking paths, and maintain robust locomotion stability.
For instance, gait libraries, which provide predefined movement references based on user commands, have been integrated into such frameworks [55]. Despite the structured approach of using gait libraries, their static nature offers limited adaptability to changing terrains, diminishing their effectiveness. A more dynamic approach involves online planning, which has shown greater adaptability and efficiency. One notable framework combines a conventional foot planner with an LL DRL policy [35], delivering targeted footsteps and directional guidance to the robot, thereby enabling responsive and varied walking commands. Moreover, HL controllers can provide additional feedback to LL policies, incorporating CoM or end-feet information, either from model-based methods or other conventional control strategies. However, this work has not yet been transferred from simulation to real-world applications. Later, a similar structure featuring an HL foot planner and an LL DRL-based policy was proposed [36]. This strategy not only achieved a successful sim-to-real transfer but also enabled the robot to navigate omnidirectionally and avoid obstacles.
A recent development has shown that focusing solely on foot placement might restrict the stability and adaptability of locomotion, particularly in complex maneuvers. A new framework integrates a model-based planner with a DRL feedback policy to enhance bipedal locomotion’s agility and versatility, displaying improved performance [58]. This system employs a residual learning architecture, where the DRL policy’s outputs are merged with the planner’s directives before being relayed to the PD controller. This integrated approach not only concerns itself with foot placement but also generates comprehensive trajectories for trunk position, orientation, and ankle yaw angle, enabling the robot to perform a wide array of locomotion skills including walking, squatting, turning, and stair climbing.
Compared to traditional model-based controllers, learned DRL policies provide a comprehensive closed-loop control strategy that does not rely on assumptions about terrain or robotic capabilities. These policies have demonstrated high efficiency in locomotion and accurate reference tracking [59]. Despite their extensive capabilities, such policies generally require short inference times, making DRL a preferred approach in scenarios where robustness is paramount or computational resources on the robot are limited. Nonetheless, these learning algorithms often face challenges in environments characterised by sparse rewards, where suitable footholds like gaps or stepping stones are infrequent [59].
Additionally, an HL planner can process critical data such as terrain variations or obstacles and generate precise target locations for feet or desired walking paths, instead of detailed terrain data, which can significantly expedite the training process [36]. This capability effectively addresses the navigational limitations observed in end-to-end frameworks. Moreover, unlike the deep planning hybrid scheme where modifications post-policy establishment can be cumbersome, this hybrid scheme offers enhanced flexibility for on-the-fly adjustments.
Despite the significant potential demonstrated by previous studies, integrating DRL-based controllers with sophisticated and complex HL planners still presents limitations compared to more integrated frameworks such as end-to-end and deep planning models. Specifically, complex HL model-based planners often require substantial computational resources to resolve problems, rely heavily on model assumptions, necessitate extensive training periods, demand large datasets for optimisation and hinder rapid deployment and iterative enhancements [59].
III-C Learned hierarchy framework
The Learned Hierarchy Framework merges a learned HL planner with an LL controller, focusing initially on refining LL policies to ensure balance and basic locomotion capabilities. Subsequently, an HL policy is developed to direct the robot towards specific targets, encapsulating a structured approach to robotic autonomy.
The genesis of this framework was within a physics engine, aimed at validating its efficiency through simulation [21]. In this setup, LL policies, informed by human motions or trajectories generated via TO, strive to track these trajectories as dictated by the HL planner while maintaining balance. An HL policy is then introduced, pre-trained with long-term task goals, to navigate the environment and identify optimal paths. This structure enabled sophisticated interactions such as guiding a biped to dribble a soccer ball towards a goal. The framework was later enhanced to include imitation learning, facilitating the replication of dynamic human-like movements within the simulation environment [22].
However, despite its structured and layered approach, which allows for the reuse of learned behaviours to achieve long-term objectives, these frameworks have predominantly been validated only in simulations. The interface designed manually between the HL planner and the LL controller sometimes leads to suboptimal behaviours, including stability issues like falling.
Expanding the application of this framework, a sim-to-real strategy for a wheeled bipedal robot was proposed, focusing the LL policy on balance and position tracking, while the HL policy enhances safety by aiding in collision avoidance and making strategic decisions based on the orientation of subgoals [38].
To fully leverage its potential, HumanPlus has been developed as a versatile framework for humanoid robots, integrating hierarchical learning, multimodal perception, and real-world imitation [56]. It employs a two-layer structure, where HIT learns from human demonstrations, trained on AMASS, and HST acts as a LL tracking controller. Additionally, binocular RGB vision input enhances perception, enabling precise loco-manipulation and dynamic locomotion tasks such as jumping, walking, folding clothes, and rearranging objects. This shadowing-based imitation learning approach improves adaptability, making it a promising framework for transferring human-like skills to robots.
Learning complex locomotion skills, particularly when incorporating navigation elements, presents a significant challenge in robotics. Decomposing these tasks into distinct locomotion and navigation components allows robots to tackle more intricate activities, such as dribbling a soccer ball [21]. As discussed in the previous section, the benefits of integrating RL-based planners with RL-based controllers have been effectively demonstrated. This combination enables the framework to adeptly manage a diverse array of environments and tasks.
Within such a framework, HL policy is optimised for strategic planning and achieving specific goals. This optimisation allows for targeted enhancements depending on the tasks at hand. Moreover, the potential for continuous improvement and adaptation through further training ensures that the system can evolve over time, improving its efficiency and effectiveness in response to changing conditions or new objectives. Despite the theoretical advantages, the practical implementation of this type of sim-to-real application for bipedal robots remains largely unexplored.
Additionally, the training process for each policy within the hierarchy demands considerable computational resources [38]. The intensive nature of this training can lead to a reliance on the simulation environment, potentially causing the system to overfit to specific scenarios and thereby fail to generalise to real-world conditions. This limitation highlights a significant hurdle that must be addressed to enhance the viability of learned hierarchy frameworks in practical applications.
Besides, for the general hierarchical framework, the transition from simulation to real-world scenarios is challenging, particularly due to the complexities involved in coordinating two layers within the control hierarchy. Ensuring seamless communication and cooperation between the HL Planner and LL Controller is essential to avoid operational discrepancies. The primary challenges include: (1) Task division complexity—while the HL Planner handles strategy and provides abstract goals, the LL Controller manages precise execution, necessitating careful coordination to avoid functional overlap and conflicts. (2) Effective communication—the HL’s abstract goals must be accurately interpreted and converted by the LL into real-time actions, especially in dynamic environments. (3) Task allocation—clear division of responsibilities between layers is crucial to prevent redundancy and ensure smooth system performance.
IV Towards a Unified Framework
Previous Section II and Section III illustrate two distinct approaches for controlling bipedal robots. Whether it is end-to-end or hierarchical, their ultimate goal is to adeptly navigate all types of terrains Both these two finally towards to the unified frameworks, can be named as Humanoid Foundation Model (HFM) and Multi-layered Adaptive Models (MLAM), which are expanded based on end-to-end, and hierarchical approaches.
In this section, we explore the conceptual design of HFM and MLAM, analyse the existing gaps and challenges in achieving these unified frameworks, and discuss recent advancements in quadruped robotics that offer valuable insights for the future development of bipedal locomotion.
IV-A Humanoid Foundation Models
Inspired by the Robot Foundation Models (RFMs) [60, 61], built on an end-to-end framework, enhance versatility by learning shared representations across different skills, HFM focus specifically on humanoid robots, which inherit advantages of RFMs but possess a higher number of DoFs and a human-like morphology, requiring more advanced motion control and adaptability. As foundation model, this structure bypasses the need for task decomposition, allowing the system to learn through large datasets and generalise across diverse locomotion tasks.
HFM with large parameters are expected to be pre-trained on large-scale robotic datasets across various platforms and later fine-tuned using DRL to adapt to specific locomotive tasks. This enables humanoid robots to continuously refine their locomotion strategies and adapt to dynamic environments through interaction. For instance, AdA [62] demonstrates a framework where an agent is first pre-trained on extensive datasets and subsequently fine-tuned with RL, allowing it to optimise policy performance and achieve near-optimal control.
Other foundation model work also provide thoughts about how to build up HFM. For example Current research on RFMs mainly about Vision-Language-Action (VLA) model on robotic arm platform, derived from VLM, VLA models directly map visual information to robot actions, enabling end-to-end decision-making without explicit intermediate representations. Recent work, such as RT-2 [63], expands upon the transformer model-based RT-1 [64], a vision-language model designed for robotic control. In RT-2, the entire model is trained using a combination of a VLM and robotic demonstrations. The actions need to be quantized as inputs and subsequently dequantized.
Challenges: The challenges of RMFs can be derived from two aspects: (1) training datasets and (2) model inference time.
  1. 1.
    Training datasets: In terms of datasets, unlike Large Language Models (LLMs) or VLMs, HMF takes multimodal inputs with heterogeneous data formats that are not inherently aligned. While LLMs process textual prompts as both input and output, and VLMs typically map images and text bidirectionally, HFM for bipedal robot must handle diverse sensory data, including vision, proprioception, motion trajectories, and object interactions, making input representation and model alignment more complex. Which lead to input representation and feature alignment challenge, which means different data modalities have different feature distributions and dimensionalities, making direct fusion difficult.
    Beside of directly making the training process difficult and high requirements of model construction, another challenge in datasets is datasets collection. Current datasets provide human motions for bipedal humanoid robot to imitate. For example, AMASS [65] consists solely of human motion capture data, while Motion-X [66] extends this by incorporating textual prompt labels corresponding to the motions. Although various references exist for robots to learn different skills, most datasets are not directly tailored for robots and require additional mapping to specific robotic platforms.
  2. 2.
    Model inference time: Bipedal robots require high-frequency action strategies to maintain stability, but real-time decision-making in HFM faces several limitations. First, RMFs using LLMs as backbones are computationally heavy, leading to slow inference due to expensive forward propagation. Second, autoregressive token generation introduces sequential dependencies, preventing parallel computation and further increasing latency. Third, onboard processors lack sufficient memory and computational power to run large models efficiently, making real-time execution challenging without model compression or hardware acceleration.
IV-B Multi-layered Adaptive Model
Unlike HFM, the MLAM follows a hierarchical framework, decomposing tasks into multiple sub-tasks. Each layer is responsible for a specific level of decision-making, with higher layers handling HL planning and lower layers executing fine-grained control. This structured approach enables the system to adapt dynamically and ultimately achieve the overall goal. Each module in the hierarchical model can be fine-tuned, replaced, or upgraded without requiring re-training of the entire model.
The discrete model is structured into three hierarchical levels: strategy, planning, and control, each progressively refining task execution. Section III has demonstrated how DRL-based policies contribute to both the HL planner and LL controller. At the highest level, the strategy layer serves as the reasoning component, enabling the robot to comprehend tasks, infer objectives, and decompose long-term goal-based tasks into structured sub-tasks, which are then passed down for execution. For example, [67] employs an LLM as a reasoning layer to interpret user commands, break them into sequential task steps, and translate them into executable code for a quadruped robot.
Challenge: In the previous Section LABEL:Sec:learned_heirarchical_framework, the challenges of the hierarchical framework were briefly discussed. However, when considering the unified framework—MLAM, additional challenges emerge that require further exploration. These challenges are:
  1. 1.
    Limited Visual Perception and Multi-Modal Integration: MLAM relies on real-time perception and decision-making, yet bipedal locomotion demands low-latency, high-precision sensory processing (e.g., vision, IMU, force sensors). Integrating these multi-modal inputs across different layers of the hierarchy remains a significant challenge. Each modality operates at different timescales and levels of abstraction, making seamless coordination and alignment difficult. Moreover, existing methods often fail to effectively capture cross-modal dependencies, limiting their applicability in dynamic scenarios.
  2. 2.
    Computational Efficiency and Real-Time Execution: MLAM must balance the quality of HL decision-making with the strict low-latency requirements of real-time control. This is especially problematic for HL reasoning models, such as LLMs, which are inherently designed for step-by-step reasoning and rely on autoregressive decoding. Generating decisions token-by-token significantly increases inference time, making these models unsuitable for tasks that demand rapid response times. These delays can lead to unstable behaviours, particularly in tasks requiring fast adaptation and precise coordination in dynamic environments.
These challenges highlight the complexity of implementing MLAM and underscore the need for innovative solutions to address these limitations in real-world applications.
IV-C Fundamental challenges
The development of unified frameworks such as HFM and MLAM aims to address the challenges of bipedal locomotion in dynamic and unstructured environments. However, achieving such frameworks requires overcoming several fundamental obstacles. Among these, balancing generalisation and precision, as well as designing streamlined frameworks capable of handling complex tasks, represent two critical challenges. These challenges are explored in the following subsections.
IV-C1 Generalisation versus precision
The challenge of balancing generalisation and precision is a fundamental obstacle in achieving the unified frameworks of HFM and MLAM. For HFM, which focuses on large-scale, versatile humanoid models, the difficulty lies in enabling the model to adapt to diverse tasks while maintaining precise control in dynamic environments. For MLAM, which emphasizes task decomposition and hierarchical control, ensuring precision across sub-tasks while maintaining a HL general policy is equally challenging.
DRL has demonstrated potential in facilitating versatile locomotion skills [31]; however, challenges such as poor velocity tracking and issues with precise control often arise. While [53] shows that deep planning combined with model-based control can achieve precise velocity tracking, and [45] illustrates successful end-to-end control for precise jumping, the creation of a policy that effectively handles both diverse tasks and precise movements remains elusive. Furthermore, [51] introduces a foot constraint policy framework, enabling precise target tracking and accurate touchdown locations. Yet, there is still no framework that comprehensively addresses the dual demands of versatility and precision in locomotion.
For HFM, these challenges are exacerbated by its reliance on diverse and large-scale datasets, which often lack the precision required for specific bipedal tasks. In contrast, MLAM struggles to achieve precision across its hierarchical layers, where HL policies must generalise across tasks while LL controllers maintain stability and accuracy in dynamic scenarios.
The difficulty in simultaneously achieving precise control and a broad range of actions in bipedal locomotion using DRL stems from several factors:
  • Complex dynamics: Bipedal locomotion involves intricate dynamics, posing a significant challenge to maintaining both dynamic motion and precision.
  • Resource intensity: Executing diverse locomotion tasks requires considerable computational power and extensive data, necessitating high-quality hardware and efficient DRL algorithms.
  • Training conflicts: Training DRL systems to achieve both precision and versatility often leads to conflicts. Designing reward functions and training policies that satisfy both criteria is inherently complex.
These challenges underscore the need for innovative solutions that balance task generalisation with control precision, bridging the gap between HFM’s large-scale adaptability and MLAM’s hierarchical precision.
IV-C2 Streamlined frameworks capable of handling complex tasks
One of the critical challenges in unifying HFM and MLAM is designing frameworks that are both simple and capable of handling complex tasks. While streamlined frameworks reduce computational overhead and potential failure points, they may lack the flexibility to adapt to diverse and dynamic environments. On the other hand, complex frameworks with multiple components, such as foot planners or terrain-specific modules [35, 36], often introduce additional computational burden and coordination challenges.
For HFM, the challenge lies in creating a model that leverages large-scale general representations without adding unnecessary layers of complexity. For MLAM, the difficulty is in coordinating hierarchical decision-making across layers without introducing redundancy or inefficiency. Achieving a balance between simplicity and capability is essential for creating robust, unified frameworks.
The advantage of maintaining simplicity in the framework lies in its ability to streamline decision-making processes, thereby reducing computational overhead and potential points of failure. To achieve an optimal framework, advancements in several key areas are essential:
  • Robust and efficient DRL algorithms: Development of algorithms that can manage high-dimensional and continuous control problems more effectively.
  • Specialised neural network architectures: Design of neural architectures tailored for specific bipedal tasks, capable of processing extensive sensory data (e.g., visual and tactile inputs), similar to the innovations presented in [52].
  • Effective reward functions: Formulation of reward functions that more accurately guide the learning process towards achieving desired behaviours and strategic outcomes.
  • Advanced computational resources: Enhancement of computational capabilities to support more intensive training and faster inference, facilitating real-time decision-making in dynamic environments.
By focusing on these developmental areas, the potential to create a unified, efficient, and less complex framework for handling complex locomotion challenges in bipedal robots is significantly increased.
Figure 4: Towards a Unified Framework: the transition from the present frameworks (end-to-end and hierarchical frameworks) to a unified framework capable of handling diverse locomotion tasks. The present state is characterised by limitations such as fine-tuning requirements, limited gait patterns, restricted terrain walking, and blind locomotion. To bridge this gap, future developments must focus on generalisation, precision, streamlined frameworks, vision-based learning, and zero-shot sim-to-real capabilities. These advancements will lead to the creation of two potential unified models: (1) a multi-layered adaptive model grounded in the hierarchical framework and (2) a large robot model built upon the end-to-end framework. The unified framework will ultimately support diverse locomotion tasks in complex environments.
V Future Directions
Section II and III introduced two mainstream learning-based frameworks, HFM and MLAM, and analysed their contributions toward achieving a unified framework.While Section IV provides a clearer understanding of the current state and the existing gap between these frameworks and the expected unified framework, further discussion is needed to outline specific directions for achieving this goal.
In this section, we outline the essential components and future research directions necessary to bridge this gap. Specifically, we identify key elements that will drive the evolution from current framework limitations to a fully integrated unified framework. A summary of these fundamental elements is provided in Fig. 4.
V-A Future directions for bipedal robots
V-A1 Loco-manipulation tasks
With the rapid development of bipedal robotics, modern bipedal robots are now equipped with upper-body structures to enhance their capabilities beyond basic locomotion tasks. These advancements enable robots to perform upper-body-assisted locomotion and loco-manipulation tasks, expanding their adaptability in complex environments.
Building on this evolution, upper-body-assisted locomotion enhances movement stability by integrating upper-lower body coordination, enabling robots to navigate challenging terrain more effectively. Beyond movement, manipulation tasks focus on precise upper-body control, requiring dexterous motion planning for object interaction and task execution. At the highest level of complexity, loco-manipulation integrates both locomotion and manipulation, demanding multi-contact interactions, dynamic balance maintenance, and real-time adaptation to unpredictable environments. These tasks necessitate full-body coordination and task-driven motion planning, making them essential benchmarks for advancing unified frameworks in humanoid robotics.
Tasks include fundamental locomotion movements like flat walking and stepping over obstacles, progressing to upper-body-assisted locomotion challenges such as climbing ladders, wall-assisted walking, jumping over obstacles, and grasping for support. Beyond locomotion, precise upper-body control is essential for manipulation tasks like turning valves, driving vehicles, using tools, and connecting fire hoses. The most complex scenarios demand full-body coordination for loco-manipulation tasks such as carrying objects while walking, opening doors while navigating, and executing disaster-response operations. These benchmarks are critical for developing robust and adaptable humanoid systems, addressing key challenges posed in DARPA’s robotics initiatives [39].
Therefore, loco-manipulation presents opportunities for humanoid robots to excel beyond purely bipedal capabilities. Few studies have addressed this integrated task; one such study [68] demonstrated a ’box transportation’ framework. This framework decomposes the task into five distinct policies, each addressing different aspects of the transportation process. However, this approach lacks efficiency and does not incorporate vision-based information, suggesting substantial room for improvement. Moreover, the challenges of managing mobile tools like scooters [69] or dynamically interacting with objects such as balls [70] introduce further complexities.
To address these challenges, a hierarchical control approach could provide a more structured and adaptable solution. Decomposing loco-manipulation into multiple layers would simplify task execution by allowing precise, modular control over individual components [53]. This method enables manual fine-tuning of task-specific controllers, ensuring better stability and adaptability to environmental variations. Additionally, a structured hierarchical approach facilitates coordinated whole-body responses, improving the efficiency of loco-manipulation interactions.
Alternatively, an end-to-end learning framework could offer a more generalised and scalable solution. By leveraging task randomisation and structured curriculum learning, bipedal robots could progressively learn loco-manipulation skills in a unified policy [48, 36, 31]. During training, such policies can also learn human-like movements from motion capture data [71, 27, 25], offering promising solutions for future integrated loco-manipulation tasks within a single, versatile policy.
V-A2 Bridging the gap from simulation to reality
Due to the large number of interactions needed for DRL algorithms, training directly on robots can lead to costly damage to hardware and the environment. Consequently, training a policy in the simulation which offer a safe and cost-effective environment for developing robotics policies, and then deploying it to the hardware illustrates significant potential and efficiency. Bridging this gap is a key step toward achieving a unified framework, as it allows for the implementation of more versatile and robust systems that can handle diverse tasks across different terrains. Furthermore, the ultimate goal of achieving zero-shot transfer, where a policy trained in simulation can be deployed without additional real-world fine-tuning, directly contributes to the unified framework’s adaptability and scalability. This capability would make the framework far more practical for handling the wide range of locomotion tasks expected from future bipedal robots.
Creating a reliable simulation environment and conducting RL training is a complex task. Advanced simulators are essential for training policies using full-order dynamics, which better represent the complex interactions robots face in the real world. Simulators like Isaac Gym [72], RoboCup3D [73], OpenAI Gym [74], and MuJoCo [75] are widely used to train policies in environments that closely mimic physical conditions. These tools help minimise the sim-to-real gap, improving the success of policy transfer to real robots. Further details on these simulators can be found in Appendix B. However, the process of sim-to-real encounters significant challenges due to the approximations and simplifications made in simulations. The process of sim-to-real faces significant challenges due to the approximations and simplifications made in simulations. Two primary issues arise: (1) Inaccurate dynamic models, contribute to discrepancies between simulation environments and real-world dynamics. They often simplify or generalise complex physical interactions, leading to a gap in performance when transitioning to reality. (2) The real environment is more diverse and complex, with varying terrain, sensor noise, and unpredictable conditions not captured in simulations. Numerous sim-to-real frameworks [43, 76, 77] have shown high efficiency and performance. Despite these advancements, a significant gap persists, exacerbated by the complexity and unpredictability of physical environments.
To bridge the sim-to-real gap, several approaches have been developed, including dynamics randomisation [44], system identification [78, 79], periodic reward composition [49], learned actuator dynamics [78, 80], regulation feedback controllers [81], and adversarial motion priors [27, 26]. In domain randomisation, there are two main methods for training policies: (1) end-to-end training using robot measurement histories, and (2) policy distillation, where an expert policy (e.g., teacher-student [49]) guides a student policy that learns from internal sensory feedback, as seen in techniques like RMA [76].
Future research in bridging the sim-to-real gap will focus on enhancing the transferability and robustness of policies across diverse environments, ultimately aiming for zero-shot transfer, where no real-world fine-tuning is required. This will not only involve improving hardware but also making simulation environments more reflective of real-world conditions. Additionally, advancing sophisticated vision-based perception and multimodal sensory integration will be crucial in further minimising the sim-to-real gap, making policies more adaptable to complex and unpredictable environments.
V-A3 Integrating foundation models
Foundation models demonstrate strong reasoning capabilities, making them valuable beyond their role in hierarchical frameworks or as strategy layers for bipedal robots as discussed in Section IV. Their integration into robotics extends beyond decision-making and offers new ways to enhance control, adaptability, and learning efficiency.
HL task planning: Foundation models, particularly LLMs and VLMs, enable robots to process multimodal inputs, enhancing task planning and execution. By integrating visual and textual data, these models enhance context-awareness, allowing robots to make more informed, adaptive decisions.
For instance, Vision-Language Model Predictive Control (VLM-PC) [82] has been effective in quadrupedal robots, integrating linguistic and visual inputs to optimise HL task planning. Similarly, in hierarchical frameworks, VLMs can act as an intermediary layer, processing sensory data before passing structured commands to lower-level controllers.
By leveraging LLMs, a unified framework could seamlessly bridge HL strategic planning with detailed task execution. Future research should focus on integrating LLMs with DRL, where LLMs guide HL strategy while DRL refines real-time execution, creating an adaptive, flexible control loop. This integration would enable robots to generalize across tasks, adapt dynamically, and make autonomous complex decisions.
Language-Assisted DRL: The design of effective reward functions is a fundamental challenge in DRL for bipedal robots. While periodic reward functions have been successfully applied to facilitate cyclic movements like walking [28], defining reward structures for non-periodic tasks, such as jumping, manipulation, and complex locomotion transitions, remains an open problem. These tasks require distinct success criteria beyond simple periodic movement cycles, making manual reward tuning inefficient and time-consuming. Additionally, optimising reward functions for multi-task learning and environmental adaptability is a major bottleneck, highlighting the need for adaptive, self-adjusting reward mechanisms.
LLMs can enhance RL by guiding policy optimisation and refining reward functions. A key challenge in DRL is reward function design, which typically requires manual tuning. LLMs offer a solution by dynamically adjusting reward signals based on linguistic descriptions of task success.
For example, Kumar et al. [83] demonstrate how LLMs refine reward functions by incorporating linguistic feedback, improving hierarchical RL frameworks. The language model processes HL instructions, which are then translated into lower-level RL tasks, simplifying complex skill acquisition and human-robot interactions. Conversely, end-to-end approaches such as AnyBipe [84] use LLMs to autonomously adjust reward functions and control strategies, reducing manual intervention. This enables more autonomous learning, allowing the model to self-optimise in diverse environments and handle complex terrain navigation.
Other implementations of foundation models in RL have been comprehensively reviewed in [60]. Future research should further explore the synergy between LLMs and DRL, with a particular focus on adaptive reward shaping and self-optimising control policies. Advancing these methodologies could pave the way for more autonomous, self-learning humanoid robots, capable of dynamically refining their decision-making processes and adapting seamlessly to diverse environments with minimal human intervention.
V-A4 Vision-based learning framework
The vision-based learning framework plays a crucial role in advancing the unified framework for bipedal robots. Integrating vision into the unified system enhances the robot’s ability to perceive and interpret its environment, making it capable of handling more complex and dynamic tasks, such as navigating rough terrains or interacting with objects in real-time. Despite the importance of vision, many current frameworks, particularly in bipedal robotics, do not fully exploit this modality [46, 53]. This can be attributed to several factors: (1) Simplified Problem: Blind locomotion reduces task complexity, focusing directly on motor control strategies rather than processing visual inputs. (2) Computation and Training Complexity: Integrating vision requires significant computational resources and robust algorithms for real-time processing, terrain identification, and object detection. (3) Advanced Control Without Vision: Previous studies have already demonstrated robust blind locomotion using proprioceptive data, such as joint angles, IMU, with vision typically enhancing precision in locomotion [52] or path planning [57].
However, recent advancements in computing power and computer vision technology—such as more efficient real-time visual processing algorithms and increasingly powerful hardware—have significantly accelerated the integration of vision-based learning into robotics. This shift has led to a surge in the popularity of vision-based approaches, allowing robots to more effectively interpret and adapt to complex environments, making them a central component in modern frameworks.
Vision-based systems are essential in human locomotion for identifying obstacles and assessing terrains, and some studies have begun to show the effectiveness of integrating vision into DRL frameworks for bipedal robots [29, 52, 36]. Building on the groundwork laid by both bipedal and quadruped robots, two promising directions have emerged:
  • Height scanner mapping: This approach, evaluated in works like [52], involves using height maps generated by scanners to inform locomotion strategies. These maps provide detailed topographical data, allowing robots to plan steps on uneven or obstructed surfaces more effectively.
  • Direct vision inputs: Directly utilising inputs from cameras, such as depth or RGB images, for real-time decision-making in RL policies [57, 85]. Although previous studies like [57] have integrated visual navigation by feeding visual information to a HL planner, the potential of direct visual inputs to RL policies has not been fully explored.
Enhancing the capability of bipedal robots to directly interpret and utilise visual data without intermediary processing can revolutionise their adaptability and efficiency in real-world scenarios. The exploration of direct vision inputs to RL-based policies represents a significant opportunity for advancing the field, potentially enabling more dynamic and responsive locomotion strategies.
V-B Insights from quadruped robots
While DRL remains an emerging technology in bipedal robotics, it has firmly established its presence in the realm of quadruped robots, another category of legged systems. The diversity of frameworks developed for quadrupeds ranges from end-to-end model-based RL designed for training in real-world scenarios, where unpredictable dynamics often prevail [86, 87], to systems that include the modelling of deformable terrain to enhance locomotion over compliant surfaces [88]. Furthermore, dynamic quadruped models facilitate highly adaptable policies [89], and sophisticated acrobatic motions are achieved through imitation learning [90].
The domain of quadruped DRL has also seen significant advancements in complex hierarchical frameworks, that integrate vision-based systems. To date, two primary versions of such hierarchical frameworks have been developed: one where a deep planning module is paired with model-based control [91] within a deep planning hybrid scheme, and another that combines model-based planning with LL DRL control [59, 92] within a feedback DRL control hybrid scheme. The latter has shown substantial efficacy; it employs a Model Predictive Control (MPC) to generate reference motions, which are then followed by a LL feedback DRL policy. Additionally, the Terrain-aware Motion Generation for Legged Robots (TAMOLS) module [93] enhances the MPC and DRL policy by providing terrain height maps for effective foothold placements across diverse environments, including those not encountered during training. However, similar hierarchical hybrid control schemes have not been thoroughly investigated within the field of bipedal locomotion.
Quadruped DRL frameworks are predominantly designed to navigate complex terrains, but efforts to extend their capabilities to other tasks are underway. These include mimicking real animals through motion capture data and imitation learning [94, 95], as well as augmenting quadrupeds with manipulation abilities. This is achieved either by adding a manipulator [96, 97] or by using the robots’ legs [98]. Notably, the research presented in [97] demonstrates that loco-manipulation tasks can be effectively managed using a single unified end-to-end framework.
Despite the progress in quadruped DRL, similar advancements have been limited for bipedal robots, particularly in loco-manipulation tasks and vision-based DRL frameworks; a combination of their inherent instability, lack of accessibility to researchers and high mechanical complexity can be attributed to the cause of this disparity between quadruped and bipedal robots. Establishing a unified framework could bridge this gap, an essential step, given the integral role of bipedal robots with upper bodies in developing fully functional humanoid systems. Moreover, the potential of hybrid frameworks that combine model-based and DRL-based methods in bipedal robots remains largely untapped.
VI Conclusion
Despite significant progress in DRL for robotics, a significant gap remains between current achievements and the development of a unified framework that can efficiently handle a wide range of complex tasks. DRL research is generally divided into two main control schemes: end-to-end and hierarchical frameworks. End-to-end frameworks have demonstrated success in handling diverse locomotion skills [31], climbing stairs [46], and navigating challenging terrains such as stepping stones [29]. Meanwhile, hierarchical frameworks provide enhanced capabilities, particularly in managing both locomotion and navigation tasks simultaneously.
Each framework brings unique strengths to the unified framework. End-to-end approaches simplify control by directly mapping inputs to outputs, while reference-based and reference-free learning methods provide the versatility needed for robots to acquire diverse locomotion skills. In contrast, hierarchical frameworks improve flexibility by structuring control into layers, allowing modular task decomposition and hybrid strategies. Together, these approaches form the basis for two potential unified frameworks: the Humanoid Robot Model from the end-to-end approach and the Multi-Layered Adaptive Model from the hierarchical approach, both designed to meet the complex demands of bipedal robot locomotion.
To bridge the existing gaps and achieve a unified framework, further development of hierarchical frameworks, particularly those equipped with advanced perception systems and integrated with model-based planners, appears promising. Such frameworks could simultaneously address issues of precision and generalisation. Moreover, the advent of LLMs presents a transformative opportunity, potentially enabling the unification of language processing and visual functionalities within robotic systems. While numerous challenges remain—ranging from the technical intricacies of framework integration to real-world application—the steady progression in control framework refinement and DRL development provides a hopeful outlook. The vision of achieving such unified framework may soon move within reach.
Appendix A Deep reinforcement learning algorithms
Figure 5: Diagram for RL algorithms catalogue
The advancement and development of RL is crucial for bipedal locomotion. Specifically, advancements in deep learning provide deep NNs serving as function approximators to empower RL with the capability to handle tasks characterised by high-dimensional and continuous spaces, by efficiently discovering condensed, low-dimensional representations of complex data. In comparison to other robots of different morphologies, such as wheeled robots, bipedal robots feature much higher DoFs and continuously interact with environments, which results in higher requirements for the DRL algorithms. Especially in the legged locomotion field, policy gradient-based algorithms are prevalent in the field of bipedal locomotion.
Designing an effective NN architecture is essential for tackling complex bipedal locomotion tasks. Multi-layer perceptrons (MLP), a fundamental NN, excel in straightforward regression tasks with lower computational resource demands. A comprehensive comparison between MLP and the memory-based NN, LSTM reveals that MLPs have an advantage in convergence speed for tasks [77]. However, LSTM, as a variant of Recurrent Neural Networks (RNN), is adept at processing data associated with time, effectively relating different states across time, and modelling key physical properties vital for periodical gaits [28] and successful sim-to-real transfer in bipedal locomotion. Additionally, Convolutional Neural Networks (CNN) specialise in spatial data processing, particularly for image-related tasks, making them highly suitable for environments where visual perception is crucial. This diverse range of NN architectures highlights the importance of selecting the appropriate model based on the specific requirements of the bipedal locomotion tasks.
Considering DRL algorithms, recent bipedal locomotion studies focus on model-free RL algorithms. Unlike model-based RL, which learns a model of the environment but may inherit biases from simulations that do not accurately reflect real-world conditions, model-free RL directly trains policies through environmental interaction without relying on an explicit environmental model. Although model-free RL requires more computational samples and resources, it can train a more robust policy allowing the robots to travel around challenging environments.
Many sophisticated model-free RL algorithms exist, which can be broadly classified into two categories: policy-based (or policy optimisation) and value-based approaches. Value-based methods e.g. Q-learning, State-Action-Reward-State-Action (SARSA), Deep Q-learning (DQN) [99] only excel in discrete action space and often struggle with high dimensional action space. Q-learning is an off-policy algorithm that directly learns the optimal Q-values, allowing it to derive the best possible actions irrespective of the current policy. SARSA, an on-policy variant, updates its Q-values based on the actual actions taken, making it robust in environments where the policy evolves during learning. DQN extends Q-learning by using deep NNs to approximate Q-values, enabling the algorithm to tackle complex state spaces, though it still faces challenges with high-dimensional action spaces due to the difficulty of accurate value estimation. In contrast, policy-based methods, such as policy gradient, can handle complex tasks but are generally less sample-efficient compared to value-based methods.
More advanced algorithms combine both policy-based methods and value-based methods. Actor-critic (AC) refers to a main idea simultaneously learning both a policy (actor) and a value function (critic), where it owns both advantages of both algorithms [100, 101]. Popular algorithms e.g. Trust region policy optimisation (TRPO) [102] and PPO based on policy-based methods, borrow ideas from AC. Moreover, there are other novel algorithms based on the AC framework, Deep Deterministic Policy Gradient (DDPG) [103], Twin Delayed Deep Deterministic Policy Gradients (TD3) [104], A2C (Advantage Actor-Critic), and A3C (Asynchronous Advantage Actor-Critic) [105], SAC (Soft Actor-Critic) [106]. Each algorithm has its strengths considering different tasks in the bipedal locomotion scenario. There are several key factors to value these algorithms such as: sample efficiency, robustness and generalisation, and implementation challenges. A comparative analysis work [107] illustrates that SAC-based algorithms excel in stability and achieve the highest scores, while their training efficiency significantly trails behind that of PPO that obtain relatively high score.
In [108], PPO demonstrates the robustness and computational economy in complex scenarios, such as bipedal locomotion, utilising fewer resources than TRPO. In terms of training time, PPO is much faster than SAC, and DDPG algorithms [107]. Besides, many works [28, 55, 44] have demonstrated its robustness and ease of implementation and combined with the flexibility to integrate with various NN architectures have made PPO the most popular choice in this field. Various work has demonstrated that PPO can conduct the exploration of walking [28], jumping [45], stair climbing [46], and stepping stones [29], which demonstrates its efficiency, robustness and generalisation.
Additionally, the DDPG algorithm integrates the Actor-Critic framework with DQN to facilitate off-policy training, further optimising sampling efficiency. In some explicit scenarios such as jumping, DDPG shows higher reward and better learning performance than PPO [30, 109]. TD3 is developed based on DDPG, and improve over the performance of the DDPG and SAC [106].
SAC improves exploration through its stochastic policy and entropy-regularised objective, which encourages the agent to maintain randomness in its actions, balancing exploration and exploitation more effectively than DDPG and TD3. Unlike PPO, which is an on-policy algorithm, SAC’s off-policy nature allows it to leverage a replay buffer, reusing past experiences to train without requiring constant interactions with the environment. This, combined with entropy maximisation, enables SAC to achieve faster convergence in complex environments where exploration is essential. SAC is also known for its stability and strong performance across a wide range of tasks [106]. While A2C offers improved efficiency and stability compared to A3C, the asynchronous update mechanism for A3C provides better capabilities for exploration and accelerating learning. Although these algorithms show their advancements, they are more challenging to apply due to algorithms’ complexity compared to PPO.
Appendix B Training Simulation Environment
The development of DRL algorithms and sim-to-real techniques has shown the requirement for high-quality simulators. Creating a reliable simulation environment and conducting RL training is difficult. Through the literature, several simulators are available, including Isaac Gym [72], robocup3D [73], Openai Gym [74], MuJoCo [75].
OpenAI developed Gym and Gymnasium providing a simplified environment for fast testing RL algorithms with the simplified bipedal model. Also, RoboCup serves as a platform for RL research and development. On the other hand, MuJoCo and Gazebo developed by DeepMind respectively are comprehensive physics engines that support robotics research and development, along with other applications needing physics simulation. Isaac Gym, a cutting-edge physics simulation and RL framework developed by NVIDIA, is a high-performance platform for training agents in intricate and lifelike environments.
One of the most crucial things is the parallelisation strategy and GPU simulations. For instance, Isaac Gym was developed to maximise the throughput of physics-based machine learning algorithms with particular emphasis on simulations that require large numbers of environment instances executing in parallel. Running the physics simulation on GPU can result in significant speedups, especially for large scenes with thousands of individual actors.